OpenStudy (anonymous):

what is the sum of the last four terms of the series 97+99+101+ . . . + 121?

OpenStudy (anonymous):

as the sum tends to 4 terms so 121+119+117+115=472:) othewise use Sn=n/2[a+l] where n=4 a=121 and l=115:)

OpenStudy (anonymous):

97+2=99 99+2=101 101+2= 103 103+2=105 105+2= 107 107+2= 109 109+2=111 111+2= 113 113+2=115 115+2=117 117+2= 119 119+2=121 115+117+119+121= 472

OpenStudy (anonymous):

ohhh, okay haha wow, i feel dumb.

OpenStudy (anonymous):

no dnt feel sometimes it just happens:)

OpenStudy (anonymous):

Dont, some things you just dont know, but now you've learned.